# Benchmarking neutrino-nucleus quasielastic scattering model predictions against a missing energy profile obtained using a monoenergetic neutrino beam

### v4

@@ IV Conclusion
The recent measurement of the missing energy distribution of 12C using a monoenergetic neutrino beam by JSNS2 was used to benchmark three nuclear models within the NEUT generator. The SF, SF, and ED-RMF nuclear models were tested alongside the impact of the NEUT FSI cascade and NucDeEx deexcitation routines. When including the effects of the NEUT FSI cascade and deexcitation and comparing to the full measurement, the SF model yields the best agreement with the measurement. Despite being more theoretically grounded and robust, the SF and ED-RMF models are excluded. Upon closer investigation, there are several reasons for this discrepancy. The ED-RMF model uses a double sided Gaussian to parametrize the $1p_{3/2}$ ground-state peak, introducing a non-physical contribution below the 1NKO threshold, which contributes significantly to the overall $\chi^{2}$. The ED-RMF model also underpredicts measurement in the bin after the $1p_{3/2}$ peak. A single-sided Gaussian would be a more natural method to model the ground state. The SF model does not have a contribution below the 1NKO threshold, but appears to overestimate the overall strength of the $1p_{3/2}$ peak. Although most neutrino scattering measurements are not sensitive to such effects due to the broadband nature of artificial neutrino fluxes, these comparisons highlight an issue in the modeling of the neutron spectral functions employed in the NEUT event generator and point out a source of discrepancy between the SF model and the JSNS2 neutron spectral function measurement.
@@ Appendix A Comparing final-state nuclear potentials
From the left sides of Figures 56 and 7, the contributions to the $\chi^{2}$ across all bins behave similar in a qualitative way. The lowest $\chi^{2}$ values are obtained by the RPWIA model, largely because of its reduced contribution from the $1p_{3/2}$ peak and increased strength in the $1s_{1/2}$ peak. Still, without the effect of the inelastic FSI cascade and nuclear de-excitations, all models fail the $p$-value test, notably due to insufficient strength at high $E_{m}$ values.

### v5

@@ IV Conclusion
The recent measurement of the missing energy distribution of 12C using a monoenergetic neutrino beam by JSNS2 was used to benchmark three nuclear models within the NEUT generator. The SF, SF, and ED-RMF nuclear models were tested alongside the impact of the NEUT FSI cascade and NucDeEx deexcitation routines.
When including the effects of the NEUT FSI cascade and deexcitation and comparing to the full measurement, the SF model yields the best agreement with the measurement. Despite being more theoretically grounded and robust, the SF and ED-RMF models are excluded. Upon closer investigation, there are several reasons for this discrepancy. The ED-RMF model uses a double sided Gaussian to parametrize the $1p_{3/2}$ ground-state peak, introducing a non-physical contribution below the 1NKO threshold, which drives the increase in the overall $\chi^{2}$. The ED-RMF model also underpredicts measurement in the bin after the $1p_{3/2}$ peak. A single-sided Gaussian would be a more natural method to model the ground state. The SF model does not have a contribution below the 1NKO threshold, but appears to overestimate the overall strength of the $1p_{3/2}$ peak and underestimate the contribution in the transition between the $1p_{3/2}$ and $1s_{1/2}$ peaks.
The inclusion of the FSI cascade has the most noticeable impact on the agreement between the measurement and the simulations, both with and without the inclusion of the 1NKO threshold. A natural extension to this work would be an in-depth study of different FSI models from different generators.
In the analysis where the 1NKO threshold is applied, the overall sensitivity of the measurement is degraded due to the exclusion of the $1p_{3/2}$ peak, which drives the discriminating power, and its strongly correlated neighbors. This leads to all models being accepted given the measurement once the FSI cascade and the NucDeEx are included. Despite this, a more detailed $\chi^{2}_{N-1}$ analysis allows to identify that $\chi^{2}$ contributions stem from different sources for different models – the SF model agreement is still driven by the transition between the $1p_{3/2}$ and $1s_{1/2}$ peaks, whereas the SF and ED-RMF models are now driven by the contribution in the tail. For ED-RMF, in particular, the larger $\chi^{2}$ contribution from the tail compared to SF could indicate a potential mis-modeling of inelastic FSI in NEUT, or indicate an incorrect treatment of SRCs.
Although most neutrino scattering measurements are not sensitive to such detailed aspects of the modeling of nuclear effects due to the broadband nature of artificial neutrino fluxes, these comparisons highlight an issue in the modeling of the neutron spectral functions employed in the NEUT event generator. The detailed statistical analysis points to several different sources of discrepancy between different NEUT models and the JSNS2 neutron spectral function measurement, and offers insight into specific ways in which each model can be modified.
@@ Appendix A Comparing final-state nuclear potentials
From the left sides of Figures 67 and 8, the contributions to the $\chi^{2}$ across all bins behave similar in a qualitative way. The lowest $\chi^{2}$ values are obtained by the RPWIA model, largely because of its reduced contribution from the $1p_{3/2}$ peak and increased strength in the $1s_{1/2}$ peak. Still, without the effect of the inelastic FSI cascade and nuclear de-excitations, all models fail the $p$-value test, notably due to insufficient strength at high $E_{m}$ values.
@@ Appendix B Prospects for improving the RMF nuclear model
Appendix C Changing the strength of the NEUT cascade
While this work uses only the NEUT generator’s cascade model, it is informative to investigate how the strength of the applied nucleon rescattering FSI impacts the obtained results. Figure 10 shows the impact of changing the FSI strength from the cascade. The $\chi^{2}$ and $p$-values are given in Table 5. The FSI strength is increased and decreased by 30% from the nominal NEUT value [^1] The 30% value is derived from Ref. [47]., which corresponds to a decrease and increase of the probability of FSI, respectively. In the case of the SF model, increasing the strength of the FSI yields a significant improvement of the $\chi^{2}$ and $p$-value, whereas decreasing it results in a slight worsening in the agreement with the measurement. For the SF model, the $\chi^{2}$ values improve (worsen) when increasing (decreasing) the FSI strength, but all variations remain rejected at the $p$-value threshold. The effect is similar for the ED-RMF model, with slightly better $\chi^{2}$ values compared to those for the SF model. In general, we find that the measurements favors increasing the strength of the FSI in all cases. In the case of the SF model, agreement is improved in both the $1p_{3/2}$ and the $1s_{1/2}$ shell peaks, as well as the transition between the $1p_{3/2}$ and the $1s_{1/2}$ peaks and the tail of the distribution. Since the ED-RMF model already overestimates the $1p_{3/2}$, the increase in FSI strength helps this agreement, but not as much as that of the SF model, in which the peak now becomes consistent with the measurement. Other comparisons of the NEUT FSI cascade to proton scattering measurements, such as those presented in Ref. [22], find a similar preference towards increasing the FSI strength in NEUT. The study presented in this section only varies a single degree of freedom within the NEUT cascade, and does not represent an exhaustive examination of the effects of FSI modeling on the agreement with the measurement. A similar analysis to the one presented in this text, but using the NuWro [26] event generator, reports that the NuWro cascade actually worsens the agreement of the ED-RMF model (with an improved missing energy profile) [24], despite having a similar description of transparency to that of NEUT. This effect is in conjunction with other different choices for nuclear effects and so the direct impact of the NuWro cascade is difficult to isolate. A dedicated study investigating the effects of other FSI cascades, such as the time-like Achilles [34] cascade, is required in order to fully characterize the impact of FSI modeling.
Figure 10: The SF, SF and ED-RMF models with changing strength of the NEUT FSI cascade. The FSI strength, determining the nucleon rescattering amount, is changed by $\pm 30$%.
Table 5: $\chi^{2}$ and $p$-values for the SF, SF{*} and ED-RMF models when changing the strength of the NEUT FSI cascade by $\pm 30$%. All samples are generated with NucDeEx off. The rows highlighted in bold are not rejected given the $p$-values. The $p$-values are rounded to two significant figures.
Model FSI strength $\chi^{2}/N_{\text{d.o.f}}$ $p$-value
SF nominal 14.11/16 0.59
SF $\mathbf{+30}$% 9.36/16 0.90
SF $\mathbf{-30}$% 16.40/16 0.43
SF nominal 44.53/16 0.00
SF +30+30% 35.06/16 0.00
SF 30-30% 53.41/16 0.00
ED-RMF nominal 43.04/16 0.00
ED-RMF +30+30% 34.35/16 0.01
ED-RMF 30-30% 48.77/16 0.00